OpenStudy (anonymous):

if x=1/2,y=-1 and z=2 what is 6x(y^2z)?

OpenStudy (linna):

have you tried pluging in the values

OpenStudy (linna):

for \(\LARGE x=\frac12 , y=-1,z=2\) you'd have.. \[\LARGE 6x(y^2z)=\] \[\LARGE 6\times \frac12[(-1)^2\times 2]\] can you solve it from here... ?

OpenStudy (anonymous):

thank you linna let me try: 6/2(+1)x2 3x2=6 is this the correct answer and the proper solution to it?

OpenStudy (linna):

that's correct. well done :)